OpenStudy (anonymous):

Write expression as a complex number in standard form and show work please!!! [(13i)/(1-2i)] The i is imaginary number. :) Thanks!

OpenStudy (anonymous):

your job is to multiply top and bottom by the conjugate of \[1-2i\] which is \[1+2i\]

OpenStudy (anonymous):

this works because it is always true that \[(a+bi)(a-bi)=a^2+b^2\] a real number

OpenStudy (anonymous):

\[\frac{13i}{1-2i}\times \frac{1+2i}{1+2i}=\frac{13i(1+2i)}{1^2+2^2}\] \[\frac{13i(1+2i)}{5}\] \[\frac{13i-23}{5}\]

OpenStudy (anonymous):

typo on last line, should be \[\frac{13i-26}{5}\] in standard form \[-\frac{26}{5}+\frac{13}{5}i\]